# Enhancing Wide-Angle Image Using Narrow-Angle View of the Same Scene

### v1

@@ 3.4 Discriminator Loss
3.4.1 Seam Consistency Loss
This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
@@ 3.4.1 Seam Consistency Loss
3.4.2 Perceptual Loss against Ground Truth
To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
@@ 4.1 Datasets
We have used two datasets to ensure robust training of our proposed model.
  • 1.
    DIV8K [19]: High-resolution images widely used for super-resolution tasks. More than 1500 photos with a maximum resolution of 8K are included in the dataset. It covers a wide range of scene contents.
  • 2.
    Landscapes Dataset [18] (LHQ 1024×1024): With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes.
@@ 4.2 Generator
As stated earlier, the generator network takes in $128\times 128\times 3$ patches and performs an initial convolution followed by a ReLU layer. The feature map is then fed into an array of $64$ residual blocks. Residual blocks assist in preventing the gradient from dying out during back propagation. This is followed by $2$ upsampling blocks, responsible for physically upscaling the patch using Pixel Shuffle layers. The upsampled patches passed through the output convolution block to adjust the channel dimension and come out as enhanced patches with a dimension of $256\times 256\times 3$.
@@ 4.3 Discriminator
We use a ResNet18 based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending. The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scaler denoting the realness of the input image.
4.4 Simulating Narrow and Wide FoVs
A high-quality narrow FoV and two wide FoV (one low-quality and one high-quality) images are required for training. High-quality wide images are used to calculate the loss, whereas low-quality wide and high-quality narrow images are used as model inputs. The original images from the dataset are treated as high-quality wide FoV images as well as the ground truth. To simulate high-quality narrow and low-quality wide-angle images from single-image datasets we have used the following techniques.
Low-quality wide FoV simulation: For the DIV8K dataset with 8K resolution, we simulated low-quality wide FoV images by first downsampling each image to 4096 × 4096 and then upsampling it back to 8K. Similarly, for the Landscapes dataset (LHQ 1024 × 1024), we downsampled each image to 512 × 512 and then upsampled it back to 1024 × 1024. This intentional degradation process mimics low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions. The description of workflow provided in this paper is in light of the Landscape dataset, that is, assuming the original image to be 1024 x 1024. The image reconstruction pipeline has been shown in the following Figure-4.
@@ 4.5 Training Details
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss. The full generator-discriminator framework was then trained for an additional 200 epochs, totaling 400K iterations. The generator consists of 16 residual blocks, each with two 3×3 convolutional layers, batch normalization, and parametric ReLU activation, allowing effective feature extraction while keeping training stable. To enhance high-frequency details, two upsampling layers using pixel shuffle progressively increase the resolution. The final 3×3 convolutional layer reconstructs the high-resolution image.
The convergence behavior of the visual quality loss for a single patch in the generator network is shown in Figure 5. It highlights the effectiveness of our proposed enhanced Gram matrix, which causes an earlier convergence by approximately 30K iterations compared to the traditional approach. This earlier stabilization suggests that our method improves the efficiency of optimization, reducing the training time required while maintaining reconstruction quality.
Figure 5: Faster convergence of visual quality loss with the enhanced Gram matrix.
To further stabilize the training and prevent mode collapse, spectral normalization was applied to the discriminator while gradient clipping controlled exploding gradients. The model was optimized using a combination of content loss (L1), perceptual loss (VGG-based), and adversarial loss, ensuring a strong balance between detail accuracy and perceptual quality.
4.6 Experiment Results
@@
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
2x Zoom 1x Zoom 1x Output
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
1x Output 0.5x Zoom 0.5x Output
5 Comparison with Prior Arts
We have evaluated our proposed methods on PSNR, SSIM and LPIPS[16] metrics over Set5[12], Set14[13] and BSD100[14] datasets. For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN[2], LatticeNet[24], ELAN[25], DLGSANet[23] and DMFFN[26] in Table 3 and Table 4.
@@ 5 Comparison with Prior Arts
We have presented our comparison with bicubic interpolation and Multiscale GAN [15] on the LPIPS metric in Table 5. It is to be noted that LPIPS values are within the range $[0,1]$, with a lower value denoting the generated image is perceptually more aligned with the ground truth.
Dataset bicubic Multiscale GAN Ours
Set5 0.5657 0.2823 0.0795
Set14 0.6114 0.3717 0.2032
BSD100 0.5726 0.4112 0.3361
Table 5: LPIPS Comparison (best marked in red)
Figure 6 presents a comparative analysis of our method against several prior super-resolution models, including bicubic interpolation, LatticeNet [24], ELAN [25], and DLGSANet [23]. The leftmost image represents the original ground truth, with the red bounding box highlighting the region of interest for focused evaluation. The comparative results illustrate variations in texture reconstruction, detail preservation, and structural fidelity across different approaches. Notably, our method, displayed in the bottom-right corner, achieves superior sharpness and consistency, effectively reducing blurring and artifacts observed in other models.
Figure 6: Visual comparison of super-resolution models, including bicubic interpolation, LatticeNet, ELAN, and DLGSANet. The red bounding box highlights the region of interest, with our method (bottom-right) achieving superior detail preservation and structural fidelity
@@ 6 Conclusion
We proposed a GAN-based approach to enhance wide-angle images by transferring fine details from a narrow-angle shot. Our method balances coverage and detail through multi-zoom image fusion, leveraging perceptual losses and an improved Gram matrix representation. The cascading lens stack further extends its application to multi-lens imaging systems.
Extensive experiments showed that our approach outperforms existing super-resolution and image synthesis methods, achieving superior PSNR, SSIM, and LPIPS scores while reducing artifacts. This opens new possibilities for computational photography, smartphone imaging, and digital zoom. Future work will focus on improving performance under extreme lighting, dynamic scenes, and real-time constraints, with potential enhancements through self-supervised learning.
During the preparation of this work, the author(s) used ChatGPT and Quillbot in order to ensure grammatical correctness of the text. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.

### v2

@@ 3.3.2 Discriminator Loss
  1. 1.
    Seam Consistency Loss This loss is aimed at eliminating the seams between the patches in the synthetic image by minimizing the perceptual difference at the boundaries of the adjacent patches. For adjacent patches $P_{i}$ and $P_{j}$, we take neighboring regions $R_{i}$ and $R_{j}$ which are $w$ pixels wide and compute the loss as,
@@
  2. 2.
    Perceptual Loss against Ground Truth To ensure the essence of the narrow FoV has been plausibly transferred to the synthetic image, the output is perceptually compared to the original wide FoV ground truth. We use the same VGG19 activation maps to find the perceptual difference $L^{(l)}_{perceptual}$ between the generated wide FoV image and the original image from the dataset.
@@ 4.1 Datasets
We have used Landscapes Dataset [28] (LHQ 1024×1024) to train the proposed model. With 90,000 high-quality images of 1024 x 1024 resolution, it includes a diverse range of urban and natural environments. These high-quality landscape images with rich textures and details challenge the method’s ability to handle complex natural scenes. We have randomly sampled 20000 + 500 images (discussed in the Training Details section) and fed into the training loop in the ascending order of pixel variance [29], to ensure “easier” images are encountered first.
Figure 4: Convergence of loss curves shown over multiple iterations
[H] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 2x Zoom 1x Zoom 1x Output [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 1x Output 0.5x Zoom 0.5x Output
4.2 Simulating Narrow and Wide FoVs
Low-quality wide FoV simulation: We downsampled each image by applying a blur kernel and adding noise following this paper [30], as well as resizing to 512 × 512, and then upsampled it (using bicubic interpolation) back to 1024 × 1024. This intentional degradation process mimics the wide image of low-quality imaging conditions while preserving the broader scene context, allowing us to test our methods under challenging visual conditions.
@@ 4.3 Generator
The generator takes in a pair of patches, each of dimension $128\times 128\times 3$. The wide FoV patch is passed through a CNN encoder with a final batchNorm layer to produce the content feature map. The narrow FoV patch initially undergoes feature extraction through the $Conv2_{2},Conv3_{2},Conv4_{2}$ layers of a VGG-19 backbone. The feature map produced by VGG-19 is enhanced by the visual properties computed from the original patch. We use 10 attention-based residual blocks, each with two $3\times 3$ convolution units, a ReLU, followed by the cross-view attention score calculation unit, outputs of which are added to the wide FoV patch. The details-infused wide FoV patch is then passed through a series of convolution layers interleaved by pixel shuffle layers to physically upscale the patch to a dimension of $256\times 256\times 3$.
@@ 4.4 Discriminator
We use a ResNet18 [31] based discriminator network that takes as input a $1024\times 1024\times 3$ image formed by stitching the enhanced patches generated by the generator using Poisson Blending [32]. The image undergoes a series of $6$ residual blocks followed by a layer of adaptive pooling and a fully connected layer outputting a single scalar denoting the realness of the input image.
4.5 Enhanced image reconstruction pipeline
Figure 5 highlights the steps undertaken to reconstruct the enhanced.
@@ 4.6 Training Details
The generator network was pre-trained for 50 epochs using content and visual losses before incorporating adversarial loss, each epoch carrying 10 sample images. The full generator-discriminator framework was then trained for an additional 2000 epochs, totaling 20.5K iterations.
The convergence behavior of the generator loss has been illustrated in Figure 4. These curves represent the first 15 epochs where the generator has been trained solo. Upper row features the raw loss curves as generated by the GAN architecture, while the lower row has been smoothened using the numpy convolve [33] method.
@@
[H] Dataset bicubic Multiscale GAN Ours Set5 0.5657 0.2823 0.0795 Set14 0.6114 0.3717 0.2032 BSD100 0.5726 0.4112 0.3361
Table 3: LPIPS Comparison (best marked in red)
Figure 6: Visual comparison of our results against the mentioned prior arts
@@
4.7 Experiment Results
5 Comparison with Prior Arts
We have evaluated our proposed methods on PSNR, SSIM and LPIPS [16] metrics over Set5 [34], Set14 [35] and BSD100 [36] datasets. For PSNR and SSIM, we have compared our results with existing technologies such as ESRGAN [7], LatticeNet [12], ELAN [13], DLGSANet [14] and DMFFN [15] in Table 4 and Table 5. We have presented our comparison with bicubic interpolation and Multiscale GAN [37] on the LPIPS metric in Table 3. Visual comparison of our method against prior super-resolution models has been presented in Figure 6.
@@ 6 Conclusion
In this work, we presented a novel approach for enhancing ultra-wide field-of-view (FoV) images by leveraging paired narrow-FoV views of the same scene. The core of our method is a patch-based cross-view attention mechanism that selectively transfers fine-grained details from narrow-FoV patches to their wide-FoV counterparts, guided by Gram matrix-based visual encoding that captures texture statistics for accurate alignment. To further refine the reconstruction, we introduce a cascading lens stack mechanism that progressively transfers details across adjacent lens images, facilitating spatial consistency and coherent integration across the full image.
We trained our model on the Landscape HQ dataset to ensure coverage of both structural diversity and fine detail. For evaluation, we used standard benchmark datasets including BSD100, Set5, and Set14. Experimental results demonstrate that our method consistently outperforms strong baselines such as ESRGAN, ELAN, DLGSANet, DMFFN, and LatticeNet in both PSNR and SSIM metrics. Qualitative comparisons further confirm the perceptual superiority of our results.
Overall, our framework successfully addresses the challenge of enhancing ultra-wide FoV images by combining cross-view attention, texture-aware encoding, and a cascading lens-based reconstruction strategy. Future extensions could explore its application to video enhancement and real-time inference scenarios.